what is the exact area of a circle with a diameter of 31 inches?
Found 2 solutions by algebrahouse.com, rfer:
Answer by algebrahouse.com(1659)

"what is the exact area of a circle with a diameter of 31 inches?"

Area of a circle = πr^2

If diameter = 31 inches
then radius = 15.5 inches

A = πr^2
= 3.14(15.5)^2 {substituted into area formula}
= 3.14(240.25) {squared 15.5}
= 754.385 in^2
www.algebrahouse.com

Answer by rfer(16322)

A=pir^2
A=3.14*15.5^2
A=754.439 sq in